# Higher-order shortest paths in hypergraphs

### v2

@@ CONCLUSION
In this work we have unveiled the higher-order organisation of shortest paths in systems with non-dyadic interactions. By investigating average path size in many real-world hypergraphs, higher-order interactions are crucial to introduce shortcuts into the system, enabling efficient communications beyond system units that would otherwise be disconnected if only pairwise segments were considered. We extended our analysis to the case of temporal hypergraphs, investigating both temporally fastest and topologically shortest paths showing how real-world connectivity cannot be reproduced by simple randomised null models.
Analysing large higher-order temporal networks by calculating all shortest/fastest paths has a very high computational cost, as it requires global information about the system. This makes the analysis particularly hard for very dense networks, such as the Copenhagen Network Study, or very large ones. In order to analyse larger systems, it would be important to develop more efficient methods, for instance based on the development of heuristics for higher-order networks that would allow the calculation of the shortest path lengths in an approximate but more efficient way, similar to what was recently done for the problem of higher-order motifs lotito2024exact . Another possible extension of the work would be to investigate the effect of the size of the time window considered in our calculations on the emerging connectivity, as previously done for the case of burstiness karsai2011small ; cencetti2021temporal . Finally, our results might be used to develop alternative attack strategies to dismantle the system by affecting shortest paths and flow. Our work investigates the connectivity properties of static and time-varying hypergraphs, contributing to a better understanding of the structural organisation of systems with higher-order interactions.
@@ METHODS
This section provides more detailed descriptions of the 25 datasets used in the above analyses.
@@
One dataset gelardi2020babboons records contact data between baboons (Baboons) and each hyperedge is a maximal clique within an specific interval of the temporal network of contacts.
@@
Higher-order reconstruction and aggregation
We use 25 freely available datasets for our analyses describing group interactions. Even though interactions involve groups, the majority of the datasets store interactions between individuals as (time, node, node)-tuples as a computational simplification even when interactions involve groups. We are here interested in investigating the different contributions of higher-order and dyadic edges to the higher-order organisation of networks. To this end, we assume that if $k$ individuals are all pairwise connected at time $t$ (i.e. they form a clique), then together they form a group of size $k$ and are promoted to a hyperedge of size $k$ in the temporal hypergraph.
Throughout our analyses, we distinguish between ‘true’ dyadic interactions and those which are an artefact of the data structures chosen to store the data digitally. We consider an edge to be a ‘pure pairwise‘ interaction when it does not form part of any larger clique, i.e. when it cannot be promoted to a group.
@@
To investigate the temporal evolution of social systems, we construct a temporal network as a sequence of static hypergraphs, each associated with a specific timestamp. In some datasets we have very fine-grained temporal information, with data logged at intervals of between 20s-5mins. This implies that many of the static snapshots have very few edges. We accordingly perform a preprocessing step and aggregate across bigger windows (e.g. grouping together snapshots of resolution 30s into groups of 5 minutes). This means that even for a temporal network, some static hypergraph snapshots may include nested hyperedges. The choice of aggregation window is determined in a data-driven manner by plotting the size of the largest connected component as a function of the size of the aggregation window and taking note of when its size saturates. We select the smallest window for which the largest connected component no longer grows.
The duration of a temporal path from node $i$ at some time $t_{1}$ to node $j$ is defined as the time elapsed between the activation time of node $i$ at $t_{1}$ and the first time $t_{2}$ that node $j$ becomes active after $t_{1}$. If multiple shortest/fastest paths exist between a pair of nodes, we select a single path uniformly and at random. When assigning sizes to the segments of a path there may be ambiguity when the segment forms part of more than one hyperedge. In this case, we select the hyperedge of minimum size and assign to the segment its size. This is inspired by the notion that smaller groups will tend to be more capable of accurately transmitting information.
@@
Because the datasets all have different time intervals, and to enable comparison of temporal path lengths across datasets, we consider only the number of timesteps for the information reported in Figures 5-7.

### v3

@@ CONCLUSION
In this work we have unveiled the higher-order organisation of shortest paths in systems with non-dyadic interactions and have shown the importance of non-dyadic ties for shortest paths. By investigating average path size in many real-world hypergraphs, we have seen that higher-order interactions are crucial to introduce shortcuts into the system, enabling efficient communications beyond system units that would otherwise be disconnected if only pairwise segments were considered. We extended our analysis to the case of temporal hypergraphs, investigating both temporally fastest and topologically shortest paths showing how real-world connectivity cannot be reproduced by simple randomised null models.
In the future it would be interesting to explore how the location and properties of hyperedges impact their contribution to shortest paths, which could be done using the notion of hypercoreness mancastroppa2024hypercores or using more sophisticated reference models such as those that are known to preserve static shortest path lengths across randomisation najajima2022randomising . Another idea to investigate whether truly higher-order links are positioned better than pairwise links could involve counting segment duplication, extending the analysis in S4 of our Supplementary.
Higher-order connectivity can be defined more generally using for example hyperedge overlap aksoy2020hypernetwork or by extending notions of flow to the higher-order case estrada2008communicability . A future avenue of exploration would be the analysis of our path size metric for networks where connectivity is defined more generally and an investigation of the changes that such definitions would induce van2008observable .
Analysing large higher-order temporal networks by calculating all shortest/fastest paths has a very high computational cost, as it requires global information about the system. This makes the analysis particularly hard for very dense networks, such as the Copenhagen Network Study, or very large ones. In order to analyse larger systems, it would be important to develop more efficient methods, for instance based on the development of heuristics for higher-order networks that would allow the calculation of shortest path lengths in an approximate but more efficient way, similar to what was recently done for the problem of higher-order motifs lotito2024exact . This would permit a clearer comparison also of the differences between shortest and fastest paths in temporal networks, an analysis that is currently limited by the computational infeasibility of calculating all temporal hyperpaths between 2 nodes. Future work could explore this avenue, considering for example the comparison of subsets of paths between nodes and future versions of themselves instead of between all nodes or other sampling schemes.
Another possible extension of the work would be to investigate the effect of the size of the time window considered in our calculations on the emerging connectivity, as previously done for the case of burstiness karsai2011small ; cencetti2021temporal . Finally, a study of the contribution of hyperlinks of different orders to the shortest paths might be used to develop alternative attack strategies to dismantle the system by affecting shortest paths.
Our work investigates the connectivity properties of static and time-varying hypergraphs, contributing to a better understanding of the structural organisation of systems with higher-order interactions.
@@ METHODS
This section provides more detailed descriptions of the 24 datasets used in the above analyses.
@@
In Section 5 of the Supplementary Materials, we additionally perform the same analyses on data describing coauthorship within geological journals across time, which is not inherently social but still a classic example of a setting in which higher-order are natural.
Higher-order reconstruction
We use 24 freely available datasets for our analyses describing group interactions. Even though interactions involve groups, the majority of the datasets store interactions between individuals as (time, node, node)-tuples as a computational simplification even when interactions involve groups. We are here interested in investigating the different contributions of higher-order and dyadic edges to the higher-order organisation of networks. To this end, we assume that if $k$ individuals are all pairwise connected at time $t$ (i.e. they form a clique), then together they form a group of size $k$ and are promoted to a hyperedge of size $k$ in the temporal hypergraph.
Throughout our analyses, we distinguish between ‘pure’ dyadic interactions and those which are an artefact of the data structures chosen to store the data digitally. We consider an edge to be a ‘pure pairwise‘ interaction when it does not form part of any larger clique, i.e. when it cannot be promoted to a group.
@@
Temporal aggregation in higher-order networks
To investigate the temporal evolution of social systems, we construct a temporal network consisting of a sequence of static hypergraphs, each associated with a specific timestamp. In some datasets we have very fine-grained temporal information with data logged at intervals of between 20s-5mins and which implies that many of the static hypergraphs are often extremely sparse. As a result, the temporal network will also be much more disconnected as nodes are active much less frequently. Following a procedure standard in the literature masuda2016guide ; holme2012temporal , we address this by performing a preprocessing step of coarse-graining where we aggregate multiple smaller snapshots into a single bigger window (e.g. grouping together snapshots of resolution 30s into groups of 5 minutes). The choice of aggregation window is determined in a data-driven manner by plotting the size of the largest connected component as a function of the size of the aggregation window and taking note of when its size saturates. We select the smallest window for which the largest connected component no longer grows.
The duration of a temporal path from node $i$ at some time $t_{1}$ to node $j$ is defined as the time elapsed between time $t_{1}$ and the first time $t_{n}$ that node $j$ appears in a snapshot (i.e. becomes active) after $t_{1}$.
In the case that multiple shortest/fastest paths exist between a given node-pair, we select a single path uniformly and at random. Whenever multiple hyperedges connect a single segment of a path, we select the hyperedge of minimum size and assign to the segment its size motivated by the intuition that smaller groups will more accurately transmit information as there are fewer places for information to flow towards estrada2008communicability . We more fully study the impact of choosing the minimum instead of alternatives in Section 3 of the Supplementary Materials.
Calculating shortest paths in static hypergraphs
Shortest paths in hypergraphs may be calculated by related them to the graph shortest path problem via a clique expansion of the hypergraph.
Specifically, a shortest path between any node pair in a hypergraph is obtained by first transforming a static hypergraph to an undirected graph by placing a pairwise edge between 2 nodes whenever they share the same hyperedge (in effect, the clique expansion of the hypergraph). We then apply Dijkstra’s shortest path algorithm cormen2001section to find the shortest paths in the pairwise graph. Finally, we assign to each pairwise link $(a,b)$ in the shortest path the associated hyperedge that containing that pair.
@@
ACKNOWLEDGEMENTS
F.B. acknowledges support from the Austrian Science Fund (FWF) through the project 10.55776/PAT1052824.